# Metric Graph Kernels via the Tropical Torelli Map

### v1

@@ B.2 Compared Graph Kernels
  • The $k$-Core Decomposition ($k$CD) Kernel. The $k$CD kernel compares graphs based on the $k$-core decomposition of a graph [20]. The $k$-core of a graph is the maximal subgraph in which all nodes have degree at least $k$. This decomposition assigns a core number to each node, indicating its topological depth or centrality within the graph. The $k$CD kernel uses the distribution of these core numbers as the basis for comparison between graphs.
@@
Table 3: Computational complexity for different graph kernels. In the table, .fit time is the time complexity to compute the features for all graphs in a dataset; .transform time is the time complexity to compute the full kernel matrix. $N$ is the number of graphs in a dataset.
@@
$k$CD $n$: # nodes; $m$: # edges;
@@
Appendix C Additional Experimental Results on Benchmark Datasets
C.1 Dataset Statistics
Table 4: Benchmark dataset statistics. In the table, $N$ is the number of graphs in a dataset; $\bar{n},\bar{m},\bar{g}$ are the average numbers of nodes, edges, and genus. In the sparsity column, “S”, “SS” and “D” are short for “Sparse”, “Semi-Sparse” and “Dense”.
Name $N$ $\bar{n}$ (std.) $\bar{m}$(std.) $\bar{g}$(std.) $\log(\bar{g}/\bar{n})$ Sparsity
AIDS 2000 15.59 (13.59) 16.20 (15.01) 1.63 (1.67) -2.26 S
BZR 405 35.75 (7.26) 38.36 (7.69) 3.61 (0.87) -2.29 S
BZR-MD 306 21.30 (4.19) 225.06 (85.83) 204.75 (81.70) 2.26 D
COIL-DEL 3900 21.54 (13.22) 54.24 (38.47) 33.70 (25.27) 0.45 SS
COX2 467 41.22 (4.03) 43.45 (4.27) 3.22 (0.42) -2.55 S
COX2-MD 303 26.28 (2.46) 335.12 (65.19) 309.84 (62.75) 2.47 D
DD 1178 284.32 (272.00) 715.66 (693.91) 432.36 (426.25) 0.42 SS
DHFR 756 42.43 (9.06) 44.54 (9.25) 3.12 (0.65) -2.61 S
DHFR-MD 393 23.87 (4.50) 283.02 (106.85) 260.15 (102.40) 2.39 D
ENZYMES 600 32.46 (14.87) 62.14 (25.50) 30.75 (13.01) -0.05 S
ER-MD 446 21.33 (6.01) 234.85 (133.47) 214.52 (127.63) 2.31 D
FRANKEN-STEIN 4337 16.83 (10.45) 17.88 (11.60) 2.07 (1.67) -2.10 S
IMDB-BINARY 1000 19.77 (10.06) 96.53 (105.60) 77.76 (98.22) 1.37 D
IMDB-MULTI 1500 13.00 (8.52) 65.94 (110.78) 53.93 (103.84) 1.42 D
MSRC-9 221 40.58 (5.27) 97.94 (15.14) 58.36 (10.18) 0.36 SS
MSRC-21 563 77.52 (12.27) 198.32 (36.81) 121.80 (25.20) 0.45 SS
MSRC-21C 209 40.28 (5.82) 96.60 (16.44) 57.33 (10.86) 0.35 SS
MUTAG 188 17.93 (4.58) 19.79 (5.68) 2.86 (1.30) -1.84 S
NCI1 4110 29.76 (13.55) 32.30 (14.93) 3.62 (1.98) -2.11 S
NCI109 4127 29.57 (13.55) 32.13 (14.96) 3.64 (1.98) -2.09 S
PROTEINS 1113 39.05 (45.76) 72.82 (84.60) 34.84 (40.56) -0.11 S
REDDIT-BINARY 2000 429.62 (554.05) 497.75 (622.99) 70.61 (85.56) -1.81 S
REDDIT-MULTI-5K 4999 508.51 (452.57) 594.87 (566.77) 90.08 (133.61) -1.73 S
C.2 Full Test Results
Table 5: Classification accuracy of 10-fold cross-validation. In the table, the accuracy is measured by percentage (%); “M” stands for out of memory. The max memory is set to be 128GB.
Name EH(std.) GS(std.) OS(std.) SP(std.) $k$CD(std.) TTE(std.) TTW(std.)
AIDS 80.00(2.44) 98.80(0.68) 57.10(37.47) 81.00(2.95) 80.10(2.07) 93.40(1.37) 90.50(1.94)
BZR 78.78(5.44) 57.54(10.70) 76.55(11.18) 78.74(3.98) 78.79(6.69) 82.74(7.67) 83.71(4.80)
BZR-MD 48.72(7.44) 51.00(8.82) 61.53(12.77) 48.71(7.13) 48.68(6.05) 48.13(8.89) 51.26(4.95)
COIL-DEL 1.00(0.33) 15.03(1.65) 14.72(2.47) 1.00(0.60) 1.01(0.51) 1.97(0.66) 3.10(1.03)
COX2 78.15(3.49) 76.23(3.85) 72.62(8.60) 78.17(4.33) 78.16(3.26) 78.17(5.14) 78.21(7.69)
COX2-MD 47.81(8.51) 48.85(6.42) 51.83(7.92) 48.20(6.53) 48.54(6.21) 52.20(6.58) 47.20(10.68)
DD 41.35(4.41) 51.69(8.87) M M M 58.66(3.35) M
DHFR 60.96(3.86) 56.10(6.61) 74.85(4.88) 61.00(4.14) 60.99(4.55) 58.86(6.19) 60.86(7.08)
DHFR-MD 67.90(5.39) 60.25(17.33) 66.13(7.37) 67.00(6.50) 67.90(5.46) 67.94(7.54) 67.95(7.40)
ENZYMES 16.00(2.60) 21.67(6.45) 22.00(6.74) 16.67(6.32) 16.83(5.40) 12.67(3.00) 18.67(6.05)
ER-MD 49.41(6.96) 46.23(10.77) 58.55(11.63) 59.41(9.95) 59.37(8.55) 59.40(5.18) 59.43(6.58)
FRANKEN-STEIN 49.76(5.78) 52.04(4.43) 50.15(5.46) 44.64(2.02) 44.64(2.74) 62.09(2.48) 62.07(2.99)
IMDB-BINARY 50.00(4.98) 47.40(3.38) 55.10(4.16) 46.60(3.61) 50.00(6.00) 47.40(2.33) 60.30(3.58)
IMDB-MULTI 33.30(4.33) 34.87(3.53) 34.53(2.93) 33.33(3.20) 33.33(4.39) 31.00(3.42) 40.40(2.35)
MSRC-9 7.23(4.15) 20.34(7.30) 14.9(9.47) 9.51(4.29) 10.00(5.68) 5.85(3.42) 5.40(3.84)
MSRC-21 3.00(2.49) 9.95(4.07) 8.52(3.16) 4.09(2.27) 5.33(1.77) 4.09(1.80) 4.61(1.17)
MSRC-21C 8.67(5.33) 18.14(8.42) 13.38(4.63) 10.00(5.81) 8.57(7.00) 11.45(8.81) 10.55(6.36)
MUTAG 66.46(7.92) 84.53(4.55) 83.51(9.38) 36.26(10.65) 31.46(7.23) 78.30(7.84) 85.18(10.43)
NCI1 49.95(1.85) 57.71(7.93) 56.11(9.56) 49.95(1.77) 49.95(2.91) 61.85(1.96) 62.06(2.03)
NCI109 49.63(1.56) 51.74(7.44) 60.67(6.15) 49.62(2.03) 49.62(1.58) 62.73(3.29) 61.76(1.99)
PROTEINS 40.43(4.23) 49.41(8.10) 44.93(4.87) 40.44(4.82) 40.43(4.08) 65.41(3.40) 71.16(3.49)
REDDIT-BINARY 50.00(3.35) M M M M 62.50(2.97) 67.35(3.83)
REDDIT-MULTI-5K 20.01(2.23) M M M M 43.89(2.08) M
Table 6: Computation time of features (.fit function in GraKeL). The wall time is measured in seconds. The edge histogram kernel is excluded.
Name Sparsity GS OS SP $k$CD TTE TTW
AIDS S 9.46 19.41 4.79 11.01 2.65 1.59
BZR S 8.29 3.87 1.87 5.19 1.08 1.21
BZR-MD D 2388.48 0.79 1.49 50.44 14.79 11.56
COIL-DEL SS 1026.84 244.50 17.03 77.47 25.78 34.13
COX2 S 10.37 5.09 3.71 7.90 1.50 1.20
COX2-MD D 9842.24 1.35 2.50 110.06 15.42 27.31
DD SS 8053.55 M M M 386.69 M
DHFR S 17.31 11.62 6.60 17.69 1.41 2.57
DHFR-MD D 4819.32 1.43 2.73 113.35 25.23 23.94
ENZYMES S 60.51 11.67 6.55 15.08 2.96 6.35
ER-MD D 5326.07 1.41 3.22 89.17 21.58 19.51
FRANKENSTEIN S 44.41 86.96 6.07 34.73 3.86 7.78
IMDB-BINARY D 8332.79 2.55 3.49 69.06 10.22 21.69
IMDB-MULTI D 12346.53 2.24 2.90 111.16 10.98 11.83
MSRC-9 SS 110.96 3.26 3.91 9.41 2.16 2.31
MSRC-21 SS 1222.51 50.79 36.71 111.15 24.60 15.29
MSRC-21C SS 104.49 3.69 1.75 8.81 3.52 3.80
MUTAG S 1.31 0.38 0.30 1.10 0.89 0.96
NCI1 S 48.30 173.60 15.51 50.22 7.34 10.86
NCI109 S 57.46 174.41 31.63 50.12 7.50 7.04
PROTEINS S 163.99 120.98 22.70 144.59 11.44 9.80
REDDIT-BINARY S M M M M 136.88 148.57
REDDIT-MULTI-5K S M M M M 374.67 M
Table 7: Computation time of full kernel matrices (.transform function in GraKeL). The wall time is measured in seconds. The edge histogram kernel is excluded.
Name Sparsity GS OS SP $k$CD TTE TTW
AIDS S 9.59 44.30 106.51 231.80 2.53 175.45
BZR S 8.57 6.49 3.63 11.81 0.73 9.70
BZR-MD D 2328.25 0.97 0.91 63.16 19.21 1410.91
COIL-DEL SS 939.85 684.63 1132.91 3631.45 30.89 5584.44
COX2 S 10.59 7.66 10.20 21.16 1.29 12.96
COX2-MD D 9801.29 1.52 1.18 132.11 23.93 3872.51
DD SS 7730.73 M M M 392.03 M
DHFR S 17.27 22.72 30.14 118.15 1.24 43.51
DHFR-MD D 5078.20 1.62 1.63 145.09 42.59 5970.10
ENZYMES S 60.50 26.08 18.49 42.26 2.65 147.04
ER-MD D 5503.19 1.48 2.78 119.88 31.80 5681.93
FRANKENSTEIN S 43.33 262.12 714.15 2198.13 3.79 1582.95
IMDB-BINARY D 8637.97 4.19 13.49 168.42 11.68 4152.41
IMDB-MULTI D 12528.71 3.99 20.94 388.92 12.59 6400.96
MSRC-9 SS 111.21 5.99 2.78 12.86 1.85 31.72
MSRC-21 SS 1004.74 109.10 68.17 257.27 25.97 1497.16
MSRC-21C SS 104.52 6.48 1.30 12.18 3.38 57.19
MUTAG S 1.52 0.77 0.52 1.83 0.35 2.64
NCI1 S 48.56 592.90 1675.78 4458.72 7.30 934.91
NCI109 S 57.46 567.12 3309.95 4490.26 7.69 994.78
PROTEINS S 162.10 222.92 171.15 979.98 11.56 648.14
REDDIT-BINARY S M M M M 133.08 9146.96
REDDIT-MULTI-5K S M M M M 359.45 M
Appendix D Additional Experimental Results on URN Datasets
D.1 Construction of URN Datasets
Table 8: Summary of URN datasets. In the table, $N$ is the number of graphs in a dataset; $R$ is global radius and $r$ is local radius, both measured in meters.
Name $N$ $R$ $r$ Cities (Coordinates) Patterns
URN2-1S 400 1000 150 Chicago: $(41.87,-87.63)$
Ufa: $(54.73,55.95)$
Gridiron
Chaotic
URN2-1M 400 2000 300 Chicago: $(41.87,-87.63)$
Ufa: $(54.73,55.95)$
GridironChaotic
URN2-2S 600 1000 150 Wuhan: $(30.52,114.35)$
London: $(51.50,-0.14)$
LinearTributary
URN2-2M 600 2000 300 Wuhan: $(30.52,114.35)$
London: $(51.50,-0.14)$
LinearTributary
URN3-S 600 1000 150 Detroit: $(42.34,-83.05)$
Omdurman: $(15.64,32.48)$
Nanchang: $(28.68,115.85)$
GridironChaoticLinear
URN3-M 750 2000 300 Seattle: $(47.62,-122.35)$
Ufa: $(54.73,55.95)$
Tianjin: $(39.12,117.17)$
GridironChaoticLinear
URN4-S 800 1000 150 Chicago: $(41.87,-87.63)$
Omdurman: $(15.64,32.48)$
Shenyang: $(41.80,123.43)$
Paris: $(48.87,2.29)$
GridironChaoticLinearTributary
URN4-M 800 2000 300 Chicago: $(41.87,-87.63)$
Kano: $(12.01,8.59)$
Shenyang: $(41.80,123.43)$
Berlin: $(52.52,13.40)$
GridironChaoticLinearTributary
D.2 Confusion Matrices
(a) URN3-S
(b) URN3-M
(c) URN4-S
(d) URN4-M

### v2

@@ E.1 Compared Graph Kernels
  • The $k$-Core Decomposition (kkCD) kernel. The kkCD kernel compares graphs based on the $k$-core decomposition of a graph (Nikolentzos et al., 2018). The $k$-core of a graph is the maximal subgraph in which all nodes have degree at least $k$. This decomposition assigns a core number to each node, indicating its topological depth or centrality within the graph. The kkCD kernel uses the distribution of these core numbers as the basis for comparison between graphs. The hyperparameter for the kkCD kernel is the maximum core number, which is typically set to be $k=3\sim 6$.
We also list the theoretical computational complexity for aforementioned graph kernels in Table 9.
E.2 SVM Settings
For the graph classification task, we use SVM with precomputed kernels. Specifically, we use the SVC implementation from scikit-learn, which is based on the soft-margin SVM (Section B.1). The solver is limited to a maximum of 10,000 iterations to ensure convergence without excessive computation time.
The evaluation of the classifier is performed using 10-fold cross-validation. Namely, the dataset is partitioned into ten disjoint subsets (folds), with each fold serving as a validation set once while the remaining nine folds form the training set. The samples are shuffled prior to splitting to reduce bias and improve the representativeness of each fold, and a fixed random seed ensures reproducibility of the splits. We use accuracy as the scoring metric.
E.3 Dataset Statistics
Table 5: Benchmark dataset statistics. In the table, NN is the number of graphs in a dataset; $\bar{n},\bar{m},\bar{g}$ are the average numbers of nodes, edges, and genus. In the sparsity column, “S”, “SS” and “D” are short for “Sparse”, “Semi-Sparse” and “Dense”.
Name NN $\bar{n}$ (std.) $\bar{m}$(std.) $\bar{g}$(std.) $\log(\bar{g}/\bar{n})$ Sparsity
AIDS 2000 15.59 (13.59) 16.20 (15.01) 1.63 (1.67) -2.26 S
BZR 405 35.75 (7.26) 38.36 (7.69) 3.61 (0.87) -2.29 S
BZR-MD 306 21.30 (4.19) 225.06 (85.83) 204.75 (81.70) 2.26 D
COIL-DEL 3900 21.54 (13.22) 54.24 (38.47) 33.70 (25.27) 0.45 $S$
COX2 467 41.22 (4.03) 43.45 (4.27) 3.22 (0.42) -2.55 S
COX2-MD 303 26.28 (2.46) 335.12 (65.19) 309.84 (62.75) 2.47 D
$D$ 1178 284.32 (272.00) 715.66 (693.91) 432.36 (426.25) 0.42 $S$
DHFR 756 42.43 (9.06) 44.54 (9.25) 3.12 (0.65) -2.61 S
DHFR-MD 393 23.87 (4.50) 283.02 (106.85) 260.15 (102.40) 2.39 D
ENZYMES 600 32.46 (14.87) 62.14 (25.50) 30.75 (13.01) -0.05 S
ER-MD 446 21.33 (6.01) 234.85 (133.47) 214.52 (127.63) 2.31 D
FRANKEN-STEIN 4337 16.83 (10.45) 17.88 (11.60) 2.07 (1.67) -2.10 S
IMDB-BINARY 1000 19.77 (10.06) 96.53 (105.60) 77.76 (98.22) 1.37 D
IMDB-MULTI 1500 13.00 (8.52) 65.94 (110.78) 53.93 (103.84) 1.42 D
MSRC-9 221 40.58 (5.27) 97.94 (15.14) 58.36 (10.18) 0.36 $S$
MSRC-21 563 77.52 (12.27) 198.32 (36.81) 121.80 (25.20) 0.45 $S$
MSRC-21C 209 40.28 (5.82) 96.60 (16.44) 57.33 (10.86) 0.35 $S$
MUTAG 188 17.93 (4.58) 19.79 (5.68) 2.86 (1.30) -1.84 S
NCI1 4110 29.76 (13.55) 32.30 (14.93) 3.62 (1.98) -2.11 S
NCI109 4127 29.57 (13.55) 32.13 (14.96) 3.64 (1.98) -2.09 S
PROTEINS 1113 39.05 (45.76) 72.82 (84.60) 34.84 (40.56) -0.11 S
REDDIT-BINARY 2000 429.62 (554.05) 497.75 (622.99) 70.61 (85.56) -1.81 S
REDDIT-MULTI-5K 4999 508.51 (452.57) 594.87 (566.77) 90.08 (133.61) -1.73 S
E.4 Full Test Results
Table 6: Classification accuracy of 10-fold cross-validation on benchmark datasets without labels. In the table, the accuracy is measured by percentage (%); “M” stands for out of memory. The max memory is set to be 128GB.
Name EH(std.) GS(std.) OS(std.) SP(std.) WL(std.) kkCD(std.) TTE(std.) TTW(std.)
AIDS 80.00(2.44) 98.80(0.68) 57.10(37.47) 81.00(2.95) 99.70(0.40) 80.10(2.07) 93.40(1.37) 90.50(1.94)
BZR 78.78(5.44) 57.54(10.70) 76.55(11.18) 78.74(3.98) 78.82(6.08) 78.79(6.69) 82.74(7.67) 83.71(4.80)
BZR-MD 48.72(7.44) 51.00(8.82) 61.53(12.77) 48.71(7.13) 64.72(6.20) 48.68(6.05) 48.13(8.89) 51.26(4.95)
COIL-DEL 1.00(0.33) 15.03(1.65) 14.72(2.47) 1.00(0.60) 5.56(0.72) 1.01(0.51) 1.97(0.66) 3.10(1.03)
COX2 78.15(3.49) 76.23(3.85) 72.62(8.60) 78.17(4.33) 78.19(5.67) 78.16(3.26) 78.17(5.14) 78.21(7.69)
COX2-MD 47.81(8.51) 48.85(6.42) 51.83(7.92) 48.20(6.53) 46.22(10.83) 48.54(6.21) 52.20(6.58) 47.20(10.68)
$D$ 41.35(4.41) 51.69(8.87) M M 49.57(9.05) M 58.66(3.35) M
DHFR 60.96(3.86) 56.10(6.61) 74.85(4.88) 61.00(4.14) 60.90(4.56) 60.99(4.55) 58.86(6.19) 60.86(7.08)
DHFR-MD 67.90(5.39) 60.25(17.33) 66.13(7.37) 67.00(6.50) 67.94(5.42) 67.90(5.46) 67.94(7.54) 67.95(7.40)
ENZYMES 16.00(2.60) 21.67(6.45) 22.00(6.74) 16.67(6.32) 17.33(3.96) 16.83(5.40) 12.67(3.00) 18.67(6.05)
ER-MD 49.41(6.96) 46.23(10.77) 58.55(11.63) 59.41(9.95) 56.04(5.30) 59.37(8.55) 59.40(5.18) 59.43(6.58)
FRANKEN-STEIN 49.76(5.78) 52.04(4.43) 50.15(5.46) 44.64(2.02) 55.66(2.19) 44.64(2.74) 62.09(2.48) 62.07(2.99)
IMDB-BINARY 50.00(4.98) 47.40(3.38) 55.10(4.16) 46.60(3.61) 48.60(4.43) 50.00(6.00) 47.40(2.33) 60.30(3.58)
IMDB-MULTI 33.30(4.33) 34.87(3.53) 34.53(2.93) 33.33(3.20) 32.20(2.27) 33.33(4.39) 31.00(3.42) 40.40(2.35)
MSRC-9 7.23(4.15) 20.34(7.30) 14.9(9.47) 9.51(4.29) 16.76(7.91) 10.00(5.68) 5.85(3.42) 5.40(3.84)
MSRC-21 3.00(2.49) 9.95(4.07) 8.52(3.16) 4.09(2.27) 6.23(3.03) 5.33(1.77) 4.09(1.80) 4.61(1.17)
MSRC-21C 8.67(5.33) 18.14(8.42) 13.38(4.63) 10.00(5.81) 19.12(9.49) 8.57(7.00) 11.45(8.81) 10.55(6.36)
MUTAG 66.46(7.92) 84.53(4.55) 83.51(9.38) 36.26(10.65) 83.48(3.84) 31.46(7.23) 78.30(7.84) 85.18(10.43)
NCI1 49.95(1.85) 57.71(7.93) 56.11(9.56) 49.95(1.77) 54.48(7.05) 49.95(2.91) 61.85(1.96) 62.06(2.03)
NCI109 49.63(1.56) 51.74(7.44) 60.67(6.15) 49.62(2.03) 54.93(5.82) 49.62(1.58) 62.73(3.29) 61.76(1.99)
PROTEINS 40.43(4.23) 49.41(8.10) 44.93(4.87) 40.44(4.82) 46.39(8.35) 40.43(4.08) 65.41(3.40) 71.16(3.49)
REDDIT-BINARY 50.00(3.35) M M M 48.70(7.03) M 62.50(2.97) 67.35(3.83)
REDDIT-MULTI-5K 20.01(2.23) M M M 28.32(6.43) M 43.89(2.08) M
Table 7: Evaluation of Computational Runtime. The wall time is measured in seconds. The edge histogram kernel and the Weisfeiler–Lehman kernel are excluded (<0.1<0.1s).
(a) Computation runtime of features (.fit function in GraKeL).
Name Sparsity GS OS SP kkCD TTE TTW
AIDS S 9.46 19.41 4.79 11.01 2.65 1.59
BZR S 8.29 3.87 1.87 5.19 1.08 1.21
BZR-MD D 2388.48 0.79 1.49 50.44 14.79 11.56
COIL-DEL $S$ 1026.84 244.50 17.03 77.47 25.78 34.13
COX2 S 10.37 5.09 3.71 7.90 1.50 1.20
COX2-MD D 9842.24 1.35 2.50 110.06 15.42 27.31
$D$ $S$ 8053.55 M M M 386.69 M
DHFR S 17.31 11.62 6.60 17.69 1.41 2.57
DHFR-MD D 4819.32 1.43 2.73 113.35 25.23 23.94
ENZYMES S 60.51 11.67 6.55 15.08 2.96 6.35
ER-MD D 5326.07 1.41 3.22 89.17 21.58 19.51
FRANKENSTEIN S 44.41 86.96 6.07 34.73 3.86 7.78
IMDB-BINARY D 8332.79 2.55 3.49 69.06 10.22 21.69
IMDB-MULTI D 12346.53 2.24 2.90 111.16 10.98 11.83
MSRC-9 $S$ 110.96 3.26 3.91 9.41 2.16 2.31
MSRC-21 $S$ 1222.51 50.79 36.71 111.15 24.60 15.29
MSRC-21C $S$ 104.49 3.69 1.75 8.81 3.52 3.80
MUTAG S 1.31 0.38 0.30 1.10 0.89 0.96
NCI1 S 48.30 173.60 15.51 50.22 7.34 10.86
NCI109 S 57.46 174.41 31.63 50.12 7.50 7.04
PROTEINS S 163.99 120.98 22.70 144.59 11.44 9.80
REDDIT-BINARY S M M M M 136.88 148.57
REDDIT-MULTI-5K S M M M M 374.67 M
(b) Computation runtime of full kernel matrices (.transform function in GraKeL).
Name Sparsity GS OS SP kkCD TTE TTW
AIDS S 9.59 44.30 106.51 231.80 2.53 175.45
BZR S 8.57 6.49 3.63 11.81 0.73 9.70
BZR-MD D 2328.25 0.97 0.91 63.16 19.21 1410.91
COIL-DEL $S$ 939.85 684.63 1132.91 3631.45 30.89 5584.44
COX2 S 10.59 7.66 10.20 21.16 1.29 12.96
COX2-MD D 9801.29 1.52 1.18 132.11 23.93 3872.51
$D$ $S$ 7730.73 M M M 392.03 M
DHFR S 17.27 22.72 30.14 118.15 1.24 43.51
DHFR-MD D 5078.20 1.62 1.63 145.09 42.59 5970.10
ENZYMES S 60.50 26.08 18.49 42.26 2.65 147.04
ER-MD D 5503.19 1.48 2.78 119.88 31.80 5681.93
FRANKENSTEIN S 43.33 262.12 714.15 2198.13 3.79 1582.95
IMDB-BINARY D 8637.97 4.19 13.49 168.42 11.68 4152.41
IMDB-MULTI D 12528.71 3.99 20.94 388.92 12.59 6400.96
MSRC-9 $S$ 111.21 5.99 2.78 12.86 1.85 31.72
MSRC-21 $S$ 1004.74 109.10 68.17 257.27 25.97 1497.16
MSRC-21C $S$ 104.52 6.48 1.30 12.18 3.38 57.19
MUTAG S 1.52 0.77 0.52 1.83 0.35 2.64
NCI1 S 48.56 592.90 1675.78 4458.72 7.30 934.91
NCI109 S 57.46 567.12 3309.95 4490.26 7.69 994.78
PROTEINS S 162.10 222.92 171.15 979.98 11.56 648.14
REDDIT-BINARY S M M M M 133.08 9146.96
REDDIT-MULTI-5K S M M M M 359.45 M
Appendix F Additional Experimental Results on URN Datasets
F.1 Construction of URN Datasets
Table 8: Summary of URN datasets. In the table, NN is the number of graphs in a dataset; $R$ is global radius and $r$ is local radius, both measured in meters.
Name NN $R$ $r$ Cities (Coordinates) Patterns
URN2-1S 400 1000 150 Chicago: (41.87,87.63)(41.87,-87.63)
Ufa: $(54.73,55.95)$
Gridiron
Chaotic
URN2-1M 400 2000 300 Chicago: (41.87,87.63)(41.87,-87.63)
Ufa: $(54.73,55.95)$
GridironChaotic
URN2-2S 600 1000 150 Wuhan: $(30.52,114.35)$
London: (51.50,0.14)(51.50,-0.14)
LinearTributary
URN2-2M 600 2000 300 Wuhan: $(30.52,114.35)$
London: (51.50,0.14)(51.50,-0.14)
LinearTributary
URN3-S 600 1000 150 Detroit: (42.34,83.05)(42.34,-83.05)
Omdurman: $(15.64,32.48)$
Nanchang: $(28.68,115.85)$
GridironChaoticLinear
URN3-M 750 2000 300 Seattle: (47.62,122.35)(47.62,-122.35)
Ufa: $(54.73,55.95)$
Tianjin: $(39.12,117.17)$
GridironChaoticLinear
URN4-S 800 1000 150 Chicago: (41.87,87.63)(41.87,-87.63)
Omdurman: $(15.64,32.48)$
Shenyang: $(41.80,123.43)$
Paris: $(48.87,2.29)$
GridironChaoticLinearTributary
URN4-M 800 2000 300 Chicago: (41.87,87.63)(41.87,-87.63)
Kano: $(12.01,8.59)$
Shenyang: $(41.80,123.43)$
Berlin: $(52.52,13.40)$
GridironChaoticLinearTributary
F.2 Confusion Matrices
(a) URN3-S
(b) URN3-M
(c) URN4-S
(d) URN4-M
Appendix G Additional Discussion on Limitations and Future Research
“De-bridging” a Graph. The tropical Torelli map ignores bridge edges as they do not belong to cycles. More generally, this leads to the question of how to “de-bridge” a graph $G$, i.e., modify $G$ to eliminate bridge edges in a reasonable way. Two potential criteria to consider would be to eliminate bridges incurring: (1) a minimal change to its topology, for example, finding a new graph $G^{\prime}$ whose persistent homology is similar to that of $G$ (Weinberger, 2011); or (2) a minimal change to its metric structure, for example, finding a $G^{\prime}$ that is close to $G$ measured under the Gromov–Hausdorff distance (Mémoli, 2012). Trees represent an extreme case of this problem, since all of their edges are bridges. Developing methods to de-bridge a graph while maintaining most of its original information is an important topic for future research.
Further Real-World Examples. Beyond road networks, we remark that our tropical Torelli framework for metric graphs are applicable to a wide range of domains where the geometry and topology of continuous structures are essential. Some examples include structural brain networks in neuroscience (Bullmore and Sporns, 2009); vascular networks (such as blood vessels) in biology (Scianna et al., 2013); and electric power grids in urban planning (Pagani and Aiello, 2013). These are interesting application areas for future work.
@@ Appendix G Additional Discussion on Limitations and Future Research
Table 9: Computational complexity for different graph kernels. In the table, .fit time is the time complexity to compute the features for all graphs in a dataset; .transform time is the time complexity to compute the full kernel matrix. NN is the number of graphs in a dataset.
@@
WL $m$: # edges; $h$: # iterations
$L$: # distinct labels
$O(N\cdot hm)$ $O(N^{2}\cdot L)$
kkCD $n$: # nodes; $m$: # edges;